Let 𝐴={1,2,3,4}. Determine the truth value of each statement:

i. βˆ€π‘₯ ∈𝐴,π‘₯+3<6

ii. βˆƒπ‘₯,π‘₯+3<6

iii. βˆƒπ‘₯,2π‘₯2 +π‘₯ =15


1
Expert's answer
2021-10-28T17:22:02-0400

i.

false: "4+3>6"


ii.

true: "1+3<6"


iii.

false:

"2\cdot1^2+1=3"

"2\cdot2^2+2=10"

"2\cdot3^2+3=21"

"2\cdot4^2+4=36"